### v2

@@ 2.2 An error estimate
For some $\kappa_{u},\kappa_{w}>1$ and ${\mu_{u}},\pi_{u},\pi_{w}>0$ suppose $b_{k},c_{k},d_{k}\geq 0$ for all $k\geq 0$ satisfy
$\kappa_{u}b_{k+1}\leq b_{k}+\pi_{u}d_{k}\quad\text{and}\quad\kappa_{w}c_{k+1}\leq c_{k}+{\mu_{u}}b_{k+1}+\pi_{w}d_{k}.$
Then, letting $\iota_{k}:=\sum_{m=1}^{k}\kappa_{u}^{-m}\kappa_{w}^{-(k+1-m)}$ (understanding that $\iota_{0}=0$), for all $k\geq 0$ and $\alpha_{u},\alpha_{w}\geq 0$, we have
(9) $R^{k+1}(\alpha_{u},\alpha_{w}):=\alpha_{u}b_{k+1}+\alpha_{w}c_{k+1}$ $\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})b_{1}+\alpha_{w}\kappa_{w}^{-k}c_{1}$
$+\sum_{j=0}^{k-1}\bigl{(}\alpha_{u}\kappa_{u}^{-(k-j)}\pi_{u}+\alpha_{w}[\iota_{k-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(k-j)}\pi_{w}]\bigr{)}d_{j+1}.$
Proof 2.9.
For $k=1$, $b_{2}\leq\kappa^{-1}_{u}b_{1}+\kappa^{-1}_{u}\pi_{u}d_{1}\text{ and }c_{2}\leq\kappa^{-1}_{w}c_{1}+\kappa^{-1}_{w}{\mu_{u}}b_{2}+\kappa^{-1}_{w}\pi_{w}d_{1}$ by assumption. Multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, then summing up, observing to cancel the two instances of $\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}b_{2}$, establishes (9).
We then take $k=n+1$, and proceed by induction, assuming (9) to hold for $k=n$. Again, $b_{n+2}\leq\kappa^{-1}_{u}b_{n+1}+\kappa^{-1}_{u}\pi_{u}d_{n+1}\text{ and }c_{n+2}\leq\kappa^{-1}_{w}c_{n+1}+\kappa^{-1}_{w}{\mu_{u}}b_{n+1}+\kappa^{-1}_{w}\pi_{w}d_{n+1}$ by assumption. As in the case $k=1$, multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, and then summing up, yields
$R^{n+2}(\alpha_{u},\alpha_{w})=\alpha_{u}b_{n+2}+\alpha_{w}\kappa^{-1}_{w}c_{n+2}$ $\leq(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})b_{n+1}+\alpha_{w}\kappa^{-1}_{w}c_{n+1}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}])d_{n+1}.$
$R^{n+2}(\alpha_{u},\alpha_{w})$ $\leq((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-n}+\alpha_{w}\kappa^{-1}_{w}\iota_{n}{\mu_{u}})b_{1}+\alpha_{w}\kappa^{-1}_{w}\kappa_{w}^{-n}c_{1}$
$+\sum_{j=0}^{n-1}\bigl{(}(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-(n-j)}\pi_{u}+\alpha_{w}\kappa^{-1}_{w}[\iota_{n-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(n-j)}\pi_{w}]\bigr{)}d_{j+1}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}])d_{n+1}$
$=(\alpha_{u}\kappa_{u}^{-(n+1)}+\alpha_{w}\mu_{u}(\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n}))b_{1}+\alpha_{w}\kappa_{w}^{-(n+1)}c_{1}$
$+\sum_{j=0}^{n}\bigl{(}\alpha_{u}\kappa_{u}^{-(n+1-j)}\pi_{u}+\alpha_{w}[(\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}){\mu_{u}}\pi_{u}+\kappa_{w}^{-(n+1-j)}\pi_{w}]\bigr{)}d_{j+1}.$
Thus we obtain (9) for $k=n+1$.
Thus, by sum formulae for arithmetic-geometric progressions [12, formula 0.113],
Lemma 2.10.
@@
Suppose Section 2.1 holds and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$. Then for any $x\in X$ and, $p\in(0,\kappa)$, we have
(13) $d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})+e_{p,k}(x),$
(15) $e_{p,k}(x)$ $:=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}p^{k}}d_{U}^{2}(u^{1},S_{u}(x^{0}))+\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}p^{k}}d_{W}^{2}(w^{1},S_{w}(x^{0}))$
$+\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}b_{X}^{2}(x^{j+1},x^{j})-\varsigma_{p}^{2}b_{X}^{2}(x,x^{k}).$
Proof 2.11.
We prove for all $s\in\mathbb{R}$ that
(16) $4sd_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))-4s^{2}\leq\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})+e_{p,k}(x).$
Maximising the left hand side with respect to $s$ then establishes (13).
By the differential transformation Section 2.1Item (iii), we have
$d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq\alpha_{u}d_{U}(u^{k+1},S_{u}(x^{k}))+\alpha_{w}d_{W}(w^{k+1},S_{w}(x^{k}))=:R^{k+1}.$
Since $\{x^{n}\}_{n=0}^{k}\subset\Omega$, the inner and adjoint tracking Section 2.1Items (i) and (ii) give
$\kappa_{u}d_{U}(u^{k+1},S_{u}(x^{k}))$ $\leq d_{U}(u^{k},S_{u}(x^{k-1}))+\pi_{u}b_{X}(x^{k},x^{k-1})\quad\text{and}\quad$
Thus, invoking Lemma 2.8 with $b_{k+1}=d_{U}(u^{k+1},S_{u}(x^{k}))$, $c_{k+1}=d_{W}(w^{k+1},S_{w}(x^{k}))$, as well as $d_{k+1}=b_{X}(x^{k+1},x^{k})$, we obtain
$R^{k+1}\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})d_{U}(u^{1},S_{u}(x^{0}))+\alpha_{w}\kappa_{w}^{-k}d_{W}(w^{1},S_{w}(x^{0}))+\sum_{j=0}^{k-1}\psi_{k-j}b_{X}(x^{j+1},x^{j}).$
Using Young’s inequality several times here, and adding the productive zero,
(17) $4sR^{k+1}\leq\frac{(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})^{2}}{\theta_{k}^{u}}d_{U}^{2}(u^{1},S_{u}(x^{0}))+\frac{(\alpha_{w}\kappa_{w}^{-k})^{2}}{\theta_{k}^{w}}d_{W}^{2}(w^{1},S_{w}(x^{0}))\\ +\sum_{j=0}^{k-1}\frac{\psi_{k-j}^{2}}{\theta_{k,j}}b_{X}^{2}(x^{j+1},x^{j})+4\left(\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}\right)s^{2}+\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})-\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})$
for any $\theta_{k}^{u},\theta_{k}^{w},\theta_{k,j}>0$. Take $\theta_{k}^{u}=p^{k}\varsigma_{p}^{-1}\pi_{u}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})$, $\theta_{k}^{w}=p^{k}\varsigma_{p}^{-1}\pi_{w}\alpha_{w}\kappa_{w}^{-k}$, and $\theta_{k,j}=\varsigma_{p}^{-1}p^{k-j}\psi_{k-j}$. Observe from (10) that $\iota_{k}\leq\kappa_{w}\iota_{k+1}$. Hence $p^{k}\iota_{k}\leq(\kappa_{w}/p)p^{k+1}\iota_{k+1}$, and further, $p^{k}\psi_{k}\leq(\overline{\kappa}/p)p^{k+1}\psi_{k+1}$, where $\overline{\kappa}/p>1$. Now
Thus rearranging (17) establishes (16). Finally, the bound in (14) on $\varsigma_{p}$ follows from Eq. 12 and $\sum_{j=0}^{\infty}(p/\kappa)^{j}=1/(1-p/\kappa)=\kappa/(\kappa-p)$.
Lemma 2.12.
Let $\{e_{p,k}(x^{k+1})\}_{k=0}^{N-1}$ be as in Lemma 2.10 with $p\geq 1$. Then
$\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})$ $\leq\frac{d_{U}^{2}(u^{1},S_{u}(x^{0}))}{\pi_{u}}\bigg{(}\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg{)}+\frac{d_{W}^{2}(w^{1},S_{w}(x^{0}))}{\pi_{w}}\bigg{(}\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg{)}.$
Proof 2.13.
We split $p^{k}e_{p,k}(x^{k+1})=:A_{k}+B_{k}+C_{k}-D_{k}$ observing (15). Now
$\sum_{k=0}^{N-1}C_{k}=\sum_{k=0}^{N-1}p^{k}\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}b_{X}^{2}(x^{j+1},x^{j})=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{k=j+1}^{N-1}\psi_{k-j}b_{X}^{2}(x^{j+1},x^{j})\\ =\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}b_{X}^{2}(x^{j+1},x^{j})\leq\sum_{j=0}^{N-2}p^{j}\varsigma_{p}^{2}b_{X}^{2}(x^{j+1},x^{j})\leq\sum_{k=0}^{N-1}D_{k}.$
Moreover, using (12) and the sum formula for geometric series, we estimate that $\sum_{k=0}^{N-1}(A_{k}+B_{k})$ is less than the right-hand side of the claim.
2.3 Smoothness of differential estimates
In this section, we derive descent- and Lipschitz-type inequalities for the approximate differential ${\widetilde{F^{\prime}}(x^{k})}$, extending these classical smoothness concepts to account for differential errors under the tracking framework.
Assume that $X$ is a normed space with $X^{*}$ its dual space, and that $F^{\prime}(x^{k})\in X^{*}$ is the Fréchet differential of $F$ at $x^{k}$. We can then produce descent inequalities and Lipschitz estimates where $\widetilde{F^{\prime}}(x^{k})$ replaces $F^{\prime}(x^{k})$. We recall that if $F^{\prime}$ is $L$-Lipschitz, it then satisfies the descent inequality
(18) $\langle F^{\prime}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{L}{2}b_{X}^{2}(x,x^{k})^{2}\quad\text{for all}\quad x,x^{k}\in X$
for the choice $b_{X}(x,x^{k})=\|x-x^{k}\|_{X}$. We will, however, allow this inequality to be satisfied by any distance expression $d_{X}$ such that that $d_{X^{*}}$ is Young to $d_{X}$. By this we mean that both expression are one-homogeneous and
In particular, we allow $d_{X}$ and $d_{X^{*}}$ to be given by (semi-)norms generated by operators. We treat the descent inequality with respect to them in the next Section 3.
Simply taking $\bar{x}=x^{k}$ in the theorem to follow, and combining with the descent inequality (18), we obtain the inexact descent inequality
(19) $\langle\widetilde{F^{\prime}}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}b_{X}^{2}(x,x^{k})-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
In convergence proofs of optimisation methods, it is frequently convenient to use the three-point descent inequality (see [7, Corollary 7.2] for the convex case, or [31, Appendix B] for the non-convex case)
(20) $\langle F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta}{2}d_{X}^{2}(x,\bar{x})-\frac{L}{2}b_{X}^{2}(x,x^{k})\quad\text{for all}\quad x,x^{k},\bar{x}\in X.$
Likewise combining the next theorem with this inequality, we obtain the inexact version
(21) $\langle\widetilde{F^{\prime}}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta-\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x})-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}b_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
In the next two theorems, while we still assume normed spaces, $F^{\prime}(x^{k})$ again no longer need be a Fréchet derivative, and can again be the arbitrary “target” of Section 2.1.
@@ 2.3 Smoothness of differential estimates
Suppose Section 2.1 holds, $X$ is a normed space, $X^{*}$ its dual with $d_{X^{*}}^{2}$ Young to $d_{X}^{2}$. Also let $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$, and pick $p\in[1,\kappa)$. Then, for $e_{p,k}$ and $\varsigma_{p}$ defined in Lemma 2.10, we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})<\infty,$
and, for any $\tilde{\gamma}>0$ and $x,\bar{x}\in X$,
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x})-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}b_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
Proof 2.15.
By the Young relationship of $d_{X}$ and $d_{X}^{*}$,
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq-\frac{1}{2\tilde{\gamma}}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x}).$
Together with (13) of Lemma 2.10, this establishes the claimed inequality. Lemma 2.12 shows the boundedness of $\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})$.
Taking $x^{*}=F^{\prime}(x)$ in the following theorem, and combining with $F^{\prime}$ being $L$-Lipschitz with respect to the distances $d_{X^{*}}$ and $d_{X}$, we can get the Lipschitz-like property with error for $\widetilde{F^{\prime}}$,
$\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x))\leq\frac{(1+\vartheta)L}{2}d_{X}^{2}(x^{k},x)+\frac{1+\vartheta^{-1}}{2}e_{\mathrm{lip},k}.$
@@
Suppose Section 2.1 holds, $d_{X^{*}}$ satisfies the triangle inequality, and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$. Then, for any $\vartheta>0$,
(22) $\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),x^{*})\leq\frac{1+\vartheta}{2}d_{X^{*}}^{2}(F^{\prime}(x^{k}),x^{*})+\frac{1+\vartheta^{-1}}{2}e_{\mathrm{lip},k}\quad\text{for all}\quad x^{*}\in X^{*},$
where
@@
Moreover, if $C>0$ is a constant independent of $N$, then
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty\quad\text{whenever}\quad\sum_{j=0}^{N-1}b_{X}^{2}(x^{j+1},x^{j})<C.$
Proof 2.17.
We apply Lemma 2.10 with $p=1$ and $x=x^{k}$. With these choices $e_{1,k}(x^{k})=e_{\mathrm{lip},k}$, and (13) reads
$d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq e_{\mathrm{lip},k}.$
With this, (22) follows after we use the triangle and Young’s inequalities to derive
$\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),x^{*})\leq\frac{1+\vartheta}{2}d_{X^{*}}^{2}(F^{\prime}(x^{k}),x^{*})+\frac{1+\vartheta^{-1}}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k})).$
For the sum of $e_{\mathrm{lip},k}$, we start with the first two terms of (23). Using (12) and the sum formula for geometric series, we bound their sum over $k\in\{0,\ldots,N-1\}$ by
$d_{U}^{2}(u^{1},S_{u}(x^{0}))\frac{\varsigma_{1}}{\pi_{u}}\bigg{(}\frac{\alpha_{u}\kappa}{\kappa-1}+\frac{\alpha_{w}\mu_{u}}{(\kappa-1)^{2}}\bigg{)}+d_{W}^{2}(w^{1},S_{w}(x^{0}))\frac{\varsigma_{1}}{\pi_{w}}\bigg{(}\frac{\alpha_{w}\kappa}{\kappa-1}\bigg{)}.$
We have $\varsigma_{1}<\infty$ by Lemma 2.10. We sum the third term of (23) over $k\in\{0,\ldots,N-1\}$, and change the order of summation to obtain
$\varsigma_{1}\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}\psi_{k-j}b_{X}^{2}(x^{j+1},x^{j})=\varsigma_{1}\sum_{j=0}^{N-2}\left(\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}\right)b_{X}^{2}(x^{j+1},x^{j})\leq\frac{\varsigma_{1}^{2}}{\overline{\kappa}}\sum_{j=0}^{N-2}b_{X}^{2}(x^{j+1},x^{j}).$
Thus $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty$ whenever $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}b_{X}^{2}(x^{j+1},x^{j})<C$.

### v3

@@ 4.4 Non-escape, quasi-Féjer monotonicity, linear convergence
Suppose Section˜4.2 holds at $\bar{x}\in X$. Then $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, and the sequence is ($p$-strongly) quasi-Féjer, i.e.,
Proof 4.11.
We first treat Section˜4.2 option Item˜(a). Fix $N\in\mathbb{N}$ and suppose $\{x^{j}\}_{j=0}^{N-1}\subset\Omega_{\bar{x}}$. Observe that $\langle\Xi x|x\rangle=0$ for all $x\in X$ by the skew-adjointness of $\Xi$. Since $0\in H(\bar{x})$, using (27) in the implicit algorithm (22), we thus get
for all $k\in\{0,\ldots,N-1\}$. By $\breve{\Lambda}\leq(1-\eta)M$ and the Pythagoras’ identity (see [35, (2.3)])
(36) $\langle M(x-z)|x-\bar{x}\rangle=\frac{1}{2}\|x-z\|_{M}^{2}+\frac{1}{2}\|x-\bar{x}\|_{M}^{2}-\frac{1}{2}\|z-\bar{x}\|_{M}^{2}\quad(x,z,\bar{x}\in X),$
we obtain
Hence $x^{N}\in\mathbb{O}_{M}(\bar{x},\delta)$. Since $x^{0}\in\Omega_{\bar{x}}$ by Section˜4.2, an inductive argument shows that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, justifying the above steps. Finally, (37) shows (35), while the claim $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ follows from (39) and $\eta>0$.
Regarding option Section˜4.2Item˜(b), arguing as above with (28) in place of (27), we get in place of (37) the estimate
Using $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$, we proceed (with $1+\gamma\geq p$) as in option Item˜(a) to establish (39), and from there onwards.
Corollary 4.12.
Suppose Section˜4.2 holds at $\bar{x}\in X$ with $p>1$ and the inequality in (29) strengthened to
Then $\|x^{N}-\bar{x}\|_{M}^{2}\to 0$ at the rate $O(p^{-N})$.
4.5 Local convergence of function values
We now proceed to function values and duality gaps. The idea of possibly assuming both Section˜4.2Item˜(a) and a relaxed version of Item˜(b), as an alternative to just the latter, is to be able to study descent at non-minimising critical points. For simplicity, we only treat sublinear convergence.
@@ 4.5 Local convergence of function values
Suppose Section˜4.2 holds at $\bar{x}\in X$ and, for a non-empty set $\hat{X}\subset X$, Eq.˜28 holds for all $\hat{x}\in\hat{X}$ with $\breve{\Lambda}=\breve{\Lambda}_{\hat{x}}\leq M$, $\gamma=\gamma_{\hat{x}}\geq 0$, and $\Omega_{\hat{x}}\supset\mathbb{O}_{M}(\bar{x},\delta)$. Then
If $\Xi=0$ and Section˜4.2 holds [^1] Since the proof of the present Theorem 4.13 shows that $x^{k}\in\mathbb{O}_{M}(\hat{x},\delta)$ for all $k\in\mathbb{N}$, to prove the required (30), it would be enough to assume that just Section 4.2Item (i) holds with $\Omega\supset\mathbb{O}_{M}(\hat{x},\delta)$., then, for all $N\in\mathbb{N}$,
(43) $F+G](x^{N})\leq\inf_{\hat{x}\in\hat{X}}[F+G](\hat{x})+\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2N}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\left(\frac{1}{N}\varepsilon_{\mathrm{desc},k}(\hat{x})+\frac{k+1}{N}\varepsilon_{\mathrm{desc},k}\right)\right).$
Proof 4.14.
Lemma˜4.10 shows for all $k\in\mathbb{N}$ that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\bigcap_{\hat{x}\in\hat{X}}\Omega_{\hat{x}}$. Hence, for any $\hat{x}\in\hat{X}$, we may follow the proof of the lemma for case Item˜(b) of Section˜4.2 to establish (40) for $\bar{x}=\hat{x}$. To reach this point, the assumption $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$ was not yet needed. Now, summing (40) over $k=0,\ldots,N-1$, we obtain
Taking the supremum over $\hat{x}\in\hat{X}$, this establishes (42).
[…]
4.6 Weak convergence
Example 4.15.
Suppose $M=A^{*}A$ for some $A\in\mathbb{L}(X;V)$ for a Hilbert space $V$. Then the $M$-semi-norm convergence clearly implies $Ax^{k}\to 0$, and consequently $Mx^{k}\to 0$. Thus $M$ is weakly admissible. In Hilbert spaces, every positive semi-definite self-adjoint operator has such a square root $A$ with $V=X$. For a convolution-based construction in the space of Radon measures, see [35, Theorem 2.4].
@@ 4.6 Weak convergence
Proof 4.17.
Lemma˜4.10 proves that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, as well as that $\sup_{N\in N}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$. The latter establishes $\|x^{k+1}-x^{k}\|_{M}\to 0$, and through admissibility for weak convergence, and (22), that $\tilde{\partial}_{k+1}=-M(x^{k+1}-x^{k})\to 0$ strongly in $X^{*}$. Moreover, Section˜4.2 yields $\|\tilde{\partial}_{k+1}-x^{*}_{k+1}\|_{X^{*}}\to 0$ for some $x^{*}_{k+1}\in H(x^{k+1})$. Consequently $x^{*}_{k+1}\to 0$. Since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$, as in the proof of Lemma˜4.10, we show the quasi-Féjer monotonicity (35) for all $\hat{x}\in\hat{X}$ and $k\in\mathbb{N}$.
  1. 1.
[…]
  2. 2.
[…]
Example 4.18.
In the setting of Section˜2 and Theorem˜2.8, the weak-*-to-strong continuity of $F^{\prime}$ can be achieved, for example, when $F(x)=\frac{1}{2}\|S(x)-b\|^{2}$ for a Lipschitz and bounded $S$ with finite-dimensional range.
Remark 4.19.
All of our theory also applies when $X$ is the dual space of a separable normed space $X_{*}$, and we replace in our definitions $X^{*}$ by the predual space $X_{*}$, that is, subdifferentials are subsets of $X_{*}$, and $M,\Lambda\in\mathbb{L}(X;X_{*})$, etc. With this change the theory applies, for example, to $X$ a space of Radon measures, as in [35]. Then Theorem˜4.16 proves the weak-* convergence.